OpenStudy (anonymous):

Alex must choose a 4-digit code for a PIN number from the digits 1 through 9, without repeating any digits. What is the probability that the selected code: a) is “1234” b) starts with a “1” c) contains a “1” in any position

OpenStudy (anonymous):

looking for help

OpenStudy (anonymous):

The total # of different combinations is 9 x 8 x 7 x 6. So for question a), you have: 1/(9 x 8 x 7 x 6) The number of combinations that start with a "1" is 1 x 8 x 7 x 6. So, for question b), you have: (8 x 7 x 6)/(9 x 8 x 7 x 6) = 1/9

OpenStudy (anonymous):

For c), the number of combinations that have a "1" anywhere in it is: 1 x 8 x 7 x 6 + 8 x 1 x 7 x 6 + 8 x 7 x 1 x 6 + 8 x 7 x 6 x 1 or 4(8 x 7 x 6) So, you would have: 4(8 x 7 x 6) / (9 x 8 x 7 x 6) = 4/9

OpenStudy (anonymous):

I really appreciate that thnx so much

OpenStudy (anonymous):

You're quite welcome! Good luck in your studies!